# UniDoorManip: Learning Universal Door Manipulation Policy Over Large-scale and Diverse Door Manipulation Environments

### 05 Mar 2024 v1

@@ 1 INTRODUCTION
In the three-stages manipulation, the first stage, handle grasping, only requires the grasp pose of the end-effector, while the latter two stages require action sequence generation. Hence we employ a policy specifically for predicting handle grasping action. Leveraging the inherent generalizability across diverse geometries provided by visual affordance [8], the policy predicts a point-level score map where a higher score indicates a greater chance of successful door manipulation, to propose the grasp action. After grasping the handle steadily, the goal of the following stages is to manipulate the handle to unlock the door and open the door. Unlike most works configuring the door with a simple manipulation mechanism, our proposed environment simulates diverse realistic door manipulation mechanisms that closely resemble real-world scenarios, e.g., lever, round, key and valve, as shown in the right part of Figure 1. Although handle manipulation and door opening share the similarity of generating action sequences, their specific action types differ significantly. Hence we train separate policies for each stage to alleviate the burden of the model, enabling a universal capability for handle manipulation and door opening. In contrast to open-loop manipulation [31], which lacks real-time adjustments, we employ a closed-loop formulation that dynamically adapts subsequent actions based on the current observation for these two stages. Furthermore, to seamlessly integrate the three separate but related universal policies for each stage into a single universal policy for the entire manipulation process, we introduce a conditioned training strategy that bridges the gap among policies.
@@ 2.1 Door Manipulation Environment and Datasets
Extensive investigations have been conducted in the field of articulated objects, encompassing various facets such as reconstructing the dynamic structure of articulated objects [1, 10, 14, 22, 18, 12], estimating 6d pose [11, 17, 15, 30, 6], comprehending the intricacies of manipulating them [13, 33, 4, 5, 21, 9, 7, 2, 24], and point-level visual affordance [20, 31, 29]. Among these areas of study, particular attention has been dedicated to the manipulation of articulated objects, with a specific emphasis on generating optimal action sequences. Previous works [27, 5] have explored the training of a single universal policy using state-based  [27] or visual-based RL [5, 7], which suffer from vast exploration space and complicated manipulation mechanisms for the door manipulation task. Otherwise, we propose a novel framework that disentangles the whole manipulation process into three stages with respective universal policies and integrates them into the whole universal policy leveraging conditioned training.
@@ 4 METHOD
As illustrated in Figure 2, we propose a novel framework that disentangles door manipulation into three distinct but related stages, each with a corresponding universal manipulation policy (Section 4.1). We leverage conditioned training to train these policies, as they have inter-dependencies, and thus they can be integrated into a unified universal policy (Section 4.2). In the first stage, we employ generalizable point-level visual affordance [8] to propose stable grasp poses (Section 4.3). In the second stage, we train a universal policy covering multiple handle manipulation mechanisms in our proposed realistic environment (Section 4.4). In the third stage, we train a policy to open doors with unlocked handles(Section 4.5). Additionally, we provide a comprehensive description of our data collection and training strategy(Section 4.6).

### 08 Mar 2024 v2

@@ 1 INTRODUCTION
In the three-stages manipulation, the first stage, handle grasping, only requires the grasp pose of the end-effector, while the latter two stages require action sequence generation. Hence we employ a policy specifically for predicting handle grasping action. Leveraging the inherent generalizability across diverse geometries provided by visual affordance [8, 35], the policy predicts a point-level score map where a higher score indicates a greater chance of successful door manipulation, to propose the grasp action. After grasping the handle steadily, the goal of the following stages is to manipulate the handle to unlock the door and open the door. Unlike most works configuring the door with a simple manipulation mechanism, our proposed environment simulates diverse realistic door manipulation mechanisms that closely resemble real-world scenarios, e.g., lever, round, key and valve, as shown in the right part of Figure 1. Although handle manipulation and door opening share the similarity of generating action sequences, their specific action types differ significantly. Hence we train separate policies for each stage to alleviate the burden of the model, enabling a universal capability for handle manipulation and door opening. In contrast to open-loop manipulation [33], which lacks real-time adjustments, we employ a closed-loop formulation that dynamically adapts subsequent actions based on the current observation for these two stages. Furthermore, to seamlessly integrate the three separate but related universal policies for each stage into a single universal policy for the entire manipulation process, we introduce a conditioned training strategy that bridges the gap among policies.
@@ 2.1 Door Manipulation Environment and Datasets
Extensive investigations have been conducted in the field of articulated objects, encompassing various facets such as reconstructing the dynamic structure of articulated objects [1, 10, 14, 23, 19, 12, 3], estimating 6d pose [11, 18, 15, 32, 6], comprehending the intricacies of manipulating them [13, 37, 4, 5, 22, 9, 7, 2, 26], and point-level visual affordance [21, 33, 31]. Among these areas of study, particular attention has been dedicated to the manipulation of articulated objects, with a specific emphasis on generating optimal action sequences. Previous works [29, 5] have explored the training of a single universal policy using state-based  [29] or visual-based RL [5, 7], which suffer from vast exploration space and complicated manipulation mechanisms for the door manipulation task. Otherwise, we propose a novel framework that disentangles the whole manipulation process into three stages with respective universal policies and integrates them into the whole universal policy leveraging conditioned training.
@@ 4 METHOD
As illustrated in Figure 2, we propose a novel framework that disentangles door manipulation into three distinct but related stages, each with a corresponding universal manipulation policy (Section 4.1). We leverage conditioned training to train these policies, as they have inter-dependencies, and thus they can be integrated into a unified universal policy (Section 4.2). In the first stage, we employ generalizable point-level visual affordance [8, 24, 38, 16] to propose stable grasp poses (Section 4.3). In the second stage, we train a universal policy covering multiple handle manipulation mechanisms in our proposed realistic environment (Section 4.4). In the third stage, we train a policy to open doors with unlocked handles(Section 4.5). Additionally, we provide a comprehensive description of our data collection and training strategy(Section 4.6).